An astronaut on the Moon dropped a tool from his space vehicle which is 5m from the ground. How long would it take for the tool to reach the ground?

1
Expert's answer
2021-11-22T10:12:10-0500

We can find time the tool takes to reach the ground from the kinematic equation:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\times5\ m}{1.62\ \dfrac{m}{s^2}}}=2.48\ s."